Prove from first principles that the function f(x,y) = xsin y + y cos L

1
Expert's answer
2021-07-05T17:19:20-0400

 f(x, y) = x sin y + y cos L


By first principle, we have


f x (x, y) = lim h"\rightarrow"0 "\dfrac{f(x + h, y) - f(x, y)}{h}"


f x (x, y) = lim h"\rightarrow"0 "\dfrac{(x + h)sin y + y cosL - (x sin y + y cos L) }{h}"


f x (x, y) = lim h"\rightarrow"0 "\dfrac{(x + h)sin y - x sin y }{h}"


f x (x, y) = lim h"\rightarrow"0 "\dfrac{ h sin y }{h}"


f x (x, y) = "sin y"






f y (x, y) = lim k"\rightarrow"0 "\dfrac{f(x , y + k) - f(x, y)}{k}"


f y (x, y) = lim k"\rightarrow"0 "\dfrac{x sin (y + k) - (y + k )cosL}{k}"


f y (x, y) = lim k"\rightarrow"0 "\dfrac{x (sin y \cos k + sink\cos y) - (y + k )cosL}{k}"


f y (x, y) = lim k"\rightarrow"0 "\dfrac{x \sin y \cos k }{k} +\dfrac{ sink \ cosy}{k} -\dfrac{ (y + k )cosL}{k}"


On substituting the limit we see that f y (x, y) does not exist.